Citric acid is triprotic (H3Cit) with acidity constants (at 25 °C) of pKa1 = 3.13, pKa2 = 4.76, and pKa3 = 6.40.

a) Write out the expressions for α_0, α_1, α_2, and α_3 as a function of Ka values and {H+}.

b) Identify the pH range in which each citrate species (H3Cit, H2Cit–, HCit2–, and Cit3–) predominates.

c) A solution is prepared by dissolving 0.10 mol of Na2HCit in enough deionized water to give a total volume of 1.00 L. Promptly after dissolving in water, the salt Na2HCit dissociates “completely” according to the following reaction: Na2HCit ⇌ 2Na+ + HCit– What is the equilibrium pH of this solution? Calculate the pH manually while assuming ideal behavior.
Answer #1

So we have a tripotic citric acid(H3Cit ), We have to write its ionic reaction and reach to its details.We have already been provided with the values of pKa1 , pKa2 and pKa3 .

H3Cit  \rightarrow H+ + H2Cit-

1-\alpha0\alpha0   \alpha0 Ka1 = [ H2Cit- ] [H+ ] / [H3Cit ] \Rightarrow\alpha0 [H+ ] / 1-\alpha0 =  Ka1

\Rightarrow\alpha0 / 1-\alpha0 =   Ka1 /  [H+ ] \Rightarrow\alpha0 =  Ka1 /  Ka1 +  [H+ ]

Similarly, H2Cit-\rightarrow HCit2- + H+ ,  \Rightarrow Ka2 =   [ HCit2- ] [H+ ] / [H2Cit- ]

\alpha0 - \alpha1   \alpha1   \alpha1 +  \alpha0\Rightarrow   [ \alpha1 +  \alpha0 ] [\alpha1 ] / [ \alpha0 - \alpha1 ] =  Ka2 substituting value of  \alpha0 from the above expression, we get  \alpha1 = ( Ka2 /  Ka2 + H+ ) ( Ka1 /  Ka1 +  [H+ ] )

Again, HCit2-\rightarrow Cit3- + H+  \Rightarrow  Ka3 =  [ Cit3- ] [H+ ] / [HCit2- ]  

\alpha1 - \alpha2\alpha2     \alpha2 + \alpha1 +  \alpha0\Rightarrow Ka3 = (\alpha2 + \alpha1 +  \alpha0 ) ( \alpha2 ) / ( \alpha1 - \alpha2 ) , Substituting the values of \alpha1 &  \alpha2 from the above expressions, we get :

\alpha2 = (Ka3 / Ka3 +  H+ )( Ka2 /  Ka2 + H+ ) ( Ka1 /  Ka1 +  [H+ ] )

c) Now we have a solution of 0.1 mole of Na2HCit in 1 litre , so it implies a concentration of 0.1 M, Now writing the ionic equation, we get

H2Cit-  \rightarrow HCit2- + H+

0.1-x x x \Rightarrow x2 / 0.1-x = Ka1 , Now, ignoring x with respect to 0.1, we get

x2 / 0.1 = Ka1 , x =  (Ka1 / 0.1)0.5 , pH = -logx = 0.5 ( -log Ka1 + log 0.1 ) = 0.5 ( pKa2 - 1) = 0.5 (4.76-1) = 1.88